OpenStudy (anonymous):

HELP!!!! ILL GIVE MEDAL Divide. Express your answer in lowest terms. 2/9 divided 1/4 = ??/??

OpenStudy (anonymous):

\[\frac{ 2 }{ 9 } \div \frac{ 1 }{ 4 } = \frac{ 8 }{ 9 }\]

OpenStudy (anonymous):

:) thanks think u can help me with some more??

OpenStudy (anonymous):

Sure. Could you close this question and post another, just tag me in the question by typing @Rachella for my assistance.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

You can give a medal by clicking "Best Response" located on the right side of the comments.